# Towards Robust Incremental Learning under Ambiguous Supervision

### v3

@@ Appendix A More Implementation Details
We utilize consistency regularization (Berthelot et al. 2019) to enhance robustness to variations in input, which assumes that a classifier should produce similar class probability for a sample and its local augmented copies. Given an image $\bm{x}_{i}$, we adopt two different data augmentation modules (Khosla et al. 2020; Cubuk et al. 2019) to obtain a weakly-augmented view $\bm{x}^{w}_{i}$ and a strongly-augmented view $\bm{x}^{s}_{i}$. During the training phase, it is imperative to ensure that the predicted outcomes of the strongly-augmented view $\bm{x}^{s}_{i}$ closely approximate the $\bm{p}_{i}$ of the weakly-augmented view, i.e., $\mathcal{L}_{\text{cr}}=\frac{1}{{|\mathcal{Y}_{t}|}}\sum\limits_{j\in{\mathcal{Y}}_{t}}{-{p_{i,j}}\log(f_{j}({\bm{x}^{s}_{i}}))}$, which is analogous to the form of $\mathcal{L}_{\text{ce}}$.
@@ Appendix B Additional Experimental Results
Results with Non-Uniform Data Generation
In practice, some labels may be more analogous to the true label than others, which makes their probability of label flipping $q$ larger than others. It means that the data generation procedure is non-uniform. Referring to (Wang et al. 2022c, b), we further test PGDR with a non-uniform data generation process on CIFAR100, with the flipping matrix 13, where each entry denotes the probability of a label being a candidate. In the current task, there is a greater likelihood of confusion between new classes and old classes. Our method still maintains a strong performance advantage. As shown in Table 4, PGDR achieves promising results, outperforming baselines in 10-blurry and 30-blurry.
@@
Method 10-blurry 30-blurry
iCaRL 53.28 56.60
BiC 55.23 58.28
WA 58.43 57.24
ANCL 51.24 59.46
PGDR 64.36 65.46
@@
We further enhance our investigation into the category perception balancing ability of PGDR by incorporating IL baselines with an additional PLL methods: PRODEN (Lv et al. 2020), a classic PLL method. As illustrated in Table 5, our method achieves a performance enhancement of 9.87% (10-blurry) and 5.97% (30-blurry) on the CIFAR100 dataset, surpassing the best baseline on IPLL with $q=0.1$. In a more challenging scenario with $q=0.2$, our approach demonstrates an improvement of 11.28% (10-blurry) and 8.98% (30-blurry) relative to the best baseline. Additionally, as presented in Table 6, our method outperforms the baselines on fine-grained datasets. This once again validates the effectiveness of PGDR in balancing category perception.
@@
In this section, we still exclude the contrastive learning and the Kullback-Leibler divergence to analyze the performance of PGDR independently.
@@
Supplementary Ablation Results on Modules.
Appendix C Pseudo-code of PGDR

### v4

@@ Appendix A More Implementation Details
We utilize consistency regularization (Berthelot et al. 2019) to enhance robustness to variations in input, which assumes that a classifier should produce similar class probability for a sample and its local augmented copies. Given an image $\bm{x}_{i}$, we adopt two different data augmentation modules (Khosla et al. 2020; Cubuk et al. 2019) to obtain a weakly-augmented view $\bm{x}^{w}_{i}$ and a strongly-augmented view $\bm{x}^{s}_{i}$. During the training phase, it is imperative to ensure that the predicted outcomes of the strongly-augmented view $\bm{x}^{s}_{i}$ closely approximate the weakly-augmented view, i.e., ${\mathcal{L}_{\text{cr}}=--\frac{1}{|\mathcal{D}^{\prime}_{t}|}\sum_{i=1}^{|\mathcal{D}^{\prime}_{t}|}\sum_{j=1}^{{|\mathcal{Y}}_{t}|}p_{ij}\log(f_{j}(\bm{x}^{s}_{i}))}$ , which involves a label $\bm{p}_{i}$ based on the prediction from the weakly-augmented view.
@@ Appendix B Additional Experimental Results
Method CIFAR100 Tiny-ImageNet
PiCO 24.26 20.56
+iCaRL 58.78 32.69
+BiC 57.13 25.48
+WA 61.50 33.59
+ER-ACE 61.35 38.23
+ANCL 56.44 23.77
+PGDR 63.26 40.41
Table 4: Accuracy comparisons on non-blurry IPLL.
Method 10-blurry 30-blurry
PiCO 31.98 45.05
+iCaRL 61.59 65.42
+BiC 65.53 66.60
+WA 66.53 67.47
+ER-ACE 65.79 67.16
+ANCL 65.30 67.37
+PGDR 67.10 68.13
@@
Method 20 tasks 25 tasks
$q=0.1$ $q=0.2$ $q=0.1$ $q=0.2$
PiCO 17.98 16.12 15.17 14.07
+iCaRL 59.75 56.70 59.75 56.65
+BiC 60.25 58.62 58.53 57.24
+WA 64.12 60.67 62.49 59.34
+ER-ACE 60.44 58.30 59.33 56.68
+ANCL 59.67 58.79 59.77 58.18
+PGDR 65.54 64.09 63.21 63.82
Table 6: Accuracy comparisons on CIFAR100 under IPLL. The best results are marked in bold and the second-best marked in underline. $q$ represents the degree of label ambiguity. IL methods are equipped with the PLL method PiCO.
Results on Non-blurry IPLL
Referring to the traditional CIL (Rebuffi et al. 2017; Zhao et al. 2020), we investigate the effectiveness of PGDR in non-blurry IPLL. We set that new samples appearing in each task are derived from new classes, i.e., 0-blurry. As show in Table 4, under the condition of a flipping probability $q$ of 0.2, our solution outperforms baselines equipped with PiCO. On the CIFAR100 and Tiny-ImageNet datasets, PGDR outperforms the best baseline by 1.76% and 2.18%. This highlights the effectiveness of our method in handling disambiguation and mitigating forgetting.
@@
Results with Non-Uniform Data Generation
In practice, some labels may be more analogous to the true label than others, which makes their probability of label flipping $q$ larger than others. It means that the data generation procedure is non-uniform. Referring to (Wang et al. 2022b, a), we further test PGDR with a non-uniform data generation process on CIFAR100, with the flipping matrix 13, where each entry denotes the probability of a label being a candidate. In the current task, there is a greater likelihood of confusion between new classes and old classes. Our method still maintains a strong performance advantage. As shown in Table 5, PGDR achieves promising results, outperforming baselines in 10-blurry and 30-blurry.
Results for Different Number of Tasks
We conduct experiments to discuss the effectiveness of our method as the number of tasks increases. We set the task quantities to be 20 and 25. As shown in Table 6, under the 10-blurry condition, our method still maintains an advantage. Specifically, in the case of 20 tasks, we outperform the best baseline by 1.42% and 3.42%. In the case of 25 tasks, our method outperforms the best baseline by 4.48%. This validates the effectiveness of our method in long-term learning.
@@
We further enhance our investigation into the category perception balancing ability of PGDR by incorporating IL baselines with an additional PLL methods: PRODEN (Lv et al. 2020), a classic PLL method. As illustrated in Table 7, our method achieves a performance enhancement of 9.87% (10-blurry) and 5.97% (30-blurry) on the CIFAR100 dataset, surpassing the best baseline in IPLL with $q=0.1$. In a more challenging scenario with $q=0.2$, our approach demonstrates an improvement of 11.28% (10-blurry) and 8.98% (30-blurry) relative to the best baseline. Additionally, as presented in Table 8, our method outperforms the baselines on fine-grained datasets. This once again validates the effectiveness of PGDR in balancing category perception.
Comparison with the SOTA IL Method
To further validate the performance advantages of our method in IPLL, we discuss the SOTA IL method, C-Flat (Bian et al. 2024). We utilize the WA version of C-Flat for our experiments. As shown in Table 9, we systematically combine C-Flat with three PLL methods to mitigate forgetting. Our findings indicate that while the combination of C-Flat with PiCO yields the best results, it still performs worse than PGDR. This reinforces the idea that the significant potential of PGDR.
Comparison with the SOTA PLL Method
We integrate IL techniques with the SOTA PLL method, TERIAL (Bao, Rui, and Zhang 2024), to further validate the effectiveness of PGDR. As shown in Table 10, our method demonstrates superior performance. However, the baselines did not perform better than those equipped with other PLL methods. We suspect this is due to TERIAL causing the model to overly focus on new categories while neglecting the retention of old knowledge, leading to accelerated forgetting of prior information.
Method $q=0.1$ $q=0.2$
10-blurry 30-blurry 10-blurry 30-blurry
PRODEN 25.26 32.95 24.12 24.81
+C-Flat 62.89 63.44 62.16 63.15
PGDR 64.40 65.54 63.83 64.58
PiCO 28.00 28.90 18.66 19.34
+C-Flat 69.45 69.92 67.49 65.80
+PGDR 69.84 71.68 68.49 70.85
PaPi 24.17 31.16 21.59 22.47
+C-Flat 63.74 64.75 60.91 60.73
+PGDR 69.09 69.64 68.47 66.84
Table 9: Accuracy comparisons on CIFAR100 under IPLL.
Method $q=0.1$ $q=0.2$
10-blurry 30-blurry 10-blurry 30-blurry
TERIAL 26.11 33.37 24.39 25.20
+iCaRL 48.10 49.89 44.39 46.21
+BiC 50.55 49.06 47.25 48.68
+WA 48.42 49.53 44.43 46.88
+ER-ACE 47.79 49.04 46.12 46.22
+ANCL 50.64 51.63 51.45 51.46
PGDR 64.40 65.54 63.83 64.58
Table 10: Accuracy comparisons on CIFAR100 under IPLL. The IL methods are combined with TERIAL.
Parameter $\beta=0.2$ $\beta=0.4$ $\beta=0.6$ $\beta=0.8$ $\beta=1.0$
Accuracy 64.08 63.46 63.64 62.73 39.82
Table 11: Performance of PGDR with label-updating parameter $\beta$ on CIFAR100.
Parameter $\alpha=0.2$ $\alpha=0.4$ $\alpha=0.6$ $\alpha=0.8$ $\alpha=1.0$
Accuracy 62.54 64.19 63.78 63.83 63.47
Table 12: Performance of PGDR with disambiguation threshold $\alpha$ on CIFAR100.
Figure 6: Performance of PGDR with varying $K$ on CIFAR100 (10-blurry).
Figure 7: Performance of PGDR with varying $N_{d}$ on CIFAR100 (10-blurry).
Figure 8: Performance of PGDR with varying $m$ on CIFAR100 (10-blurry).
Figure 9: Performance comparison of PGDR, PGDR w/o CR, and PGDR w/o KD on CIFAR100 (10-blurry).
@@
Ablation Disambiguation Memory Replay 10-blurry 30-blurry
PGDR 68.49 70.85
PGDR w MP Model Prediction 67.90 68.36
PGDR w PP Prototype Prediction 30.73 27.05
PGDR w/o Memory $\checkmark$ $\times$ 26.65 26.64
PGDR w Random Random 63.03 65.99
PGDR w Distance Prototype Distance 67.15 68.34
Table 14: Ablation study of PGDR with contrastive learning on CIFAR100 with $q=0.2$ at IPLL 10 tasks.
@@
In this section, we still exclude the contrastive learning and the Kullback-Leibler divergence to analyze the performance of PGDR independently. Additionally, we conduct an analysis of PGDR’s performance when incorporating contrastive learning.
Effects of the label-updating parameter $\beta$.
We analyze the parameter $\beta$ used in the label updating based on the prediction of the model. Table 11 shows the performance of PGDR with varying $\beta$ on CIFAR100. When the $\beta$ is too high, pseudo-labels are primarily composed of disambiguated labels obtained in the initial stage. It becomes challenging to further disambiguate pseudo-labels using model outputs, leading to a decline in performance. Meanwhile, within a broader range of $\beta$ variations, the changes in the model’s classification performance are relatively small, indicating the robustness of our approach.
Effects of disambiguation threshold $\alpha$.
We analyze the confidence threshold $\alpha$ used in the label disambiguation module to distinguish between new and old classes. Table 12 shows the performance of PGDR with varying $\alpha$ on CIFAR100. It can be observed that as the threshold increases, there is a slight fluctuation in accuracy. This reflects the robust characteristics of the model towards thresholds $\alpha$ in our solution, demonstrating its strong adaptability to dynamic environments.
Effects of selection parameters $K$ and $N_{d}$.
We analyze the number $K$ of nearest neighbor samples and the storage quantity $N_{d}$ of diverse samples within the data replay module. As $N_{d}$ is determined by the product of the single-class sample storage limit and the preset ratio, we analyze the preset ratio corresponding to $N_{d}$. In Figure 7, it is evident that inadequate selection of neighboring samples results in difficulty representing local samples, consequently leading to the ineffectiveness of the diversity strategy. Conversely, when the number of selected neighboring samples is excessively high, it results in a limited diversity of filtered samples, hindering their effectiveness. In Figure 7, it can be observed that a limited quantity of diverse samples in storage is unfavorable for adequately covering the sample space, resulting in a decrease in accuracy. When storing an excessive number of diverse samples, it reduces the proportion of representative samples in storage, impacting the availability of stored data.
The impact of storage limitation $m$.
To analyze the impact of storage limits on mitigating the effects of category-aware perception balance, we set the storage limit $m$ at 500, 1000, 2000, 3000, and 4000. The experiment is conducted on 10-blurry in CIFAR100 with a flipping probability $q$=0.2. As shown in Figure 9, with the arrival of new data, when storage capacity is low, there is a catastrophic decline in performance, clearly demonstrating that increasing the storage capacity effectively alleviates catastrophic forgetting. It assists the model in recognizing old class patterns and mitigates bias towards new classes.
Figure 12: Ablation experiment of the contrastive learning module on CIFAR100. The sets include 10-blurry ($q=0.1$ and $q=0.2$) and 30-blurry ($q=0.1$ and $q=0.2$).
The impact of $\mathcal{L}_{\text{kd}}$ and $\mathcal{L}_{\text{cr}}$.
We analyze the contributions of consistency regularization (CR) and knowledge distillation (KD) within our framework. Specifically, we compare PGDR with two variants: 1) PGDR w/o CR removes consistency regularization; 2) PGDR w/o KD removes knowledge distillation. In Figure 9, it is evident that PGDR w/o CR performs worse than both PGDR and PGDR w/o KD in each task. This reflects the high reliability of our disambiguation strategy, which can provide accurate information for consistency regularization to facilitate the learning of ambiguous samples. The effectiveness of PGDR w/o KD did not experience a significant decline with the arrival of new data, which demonstrates the excellent performance of our replay strategy in mitigating forgetting.
Supplementary ablation results on modules.
@@
Ablation of PGDR with PiCO’s contrastive learning.
We conduct ablation experiments to verify the effectiveness of each component of PGDR with contrastive learning. Specifically, for the label disambiguation module, we choose the variants: 1) PGDR w MP and 2) PGDR w PP. To assess the effectiveness of the memory update strategy, we compare PGDR with three variants: 1) PGDR w/o Memory, 2) PGDR w Random and 3) PGDR w Distance. As shown in Table 14, PGDR outperforms other variants, demonstrating the superiority of our method. Specifically, in the disambiguation module’s ablation experiment, rhe effectiveness of PGDR w PP is significantly reduced. This is because the addition of contrastive learning in our method intensifies the accumulation of errors in frequent prototype updates. In the experiment of memory update strategies, we outperform three variants, demonstrating that our memory update strategy remains effective even with the addition of contrastive learning. In addition, the performance of our method remains basically consistent before and after the removal of the contrastive learning module in the Figure 12. This indicates that the CL module of PiCO is not the fundamental factor affecting PGDR performance, thus providing indirect validation of the feasibility of the proposed architecture.
Figure 13: Performance of the prototype-based separation mechanism on CIFAR100.
Method $q=0.1$ $q=0.2$
10-blurry 30-blurry 10-blurry 30-blurry
PGDR 69.84 71.68 68.49 70.85
PGDR with Conf 66.79 65.89 65.66 57.71
PGDR with Entropy 66.67 65.73 66.02 56.27
Table 15: Ablation experiment of the prototype-based separation mechanism on CIFAR100.
CIFAR100 Tiny-ImageNet
iCaRL 7.7 32.6
BiC 9.0 35.3
WA 7.8 33.6
ER-ACE 7.6 31.6
ANCL 9.8 39.5
PGDR 8.5 34.2
Table 16: Running time (seconds/epoch) of each comparing approach on CIFAR100 and Tiny-ImageNet.
The effect of the old/new data separation mechanism.
The mean accuracy of the separation mechanism is 92.32% on CIFAR100. Specifically, 98.90% of new-class samples are correctly detected as novel, and 86.49% of true old-class samples are accurately identified. As shown in Figure 13, accuracy declines with more tasks, with the most rapid decrease occurring in old-class samples due to catastrophic forgetting. We also include comparisons with separation strategies such as calculating model confidence or uncertainty, to validate the effectiveness of prototype-based separation. As shown in Table 15, our results demonstrate that as the degree of blur increases, both the confidence-based separation strategy and the entropy-based uncertainty method lead to a performance decline. We attribute this phenomenon to catastrophic forgetting, which causes the model to exhibit low confidence and high uncertainty for both the forgotten old knowledge and the previously unseen new knowledge, thereby significantly reducing the separation efficacy.
Appendix C Complexity Analysis
Let $P$, $B$, $D$, and $K$ denote the dimension of the input sample, the batch size, the dimensionality of the feature, and the number of classes. Let $H$ denote the hidden dimensionalities of the model. Our time complexity is $O(B(PH+HD+DK))$. Below we show the running time of PGDR and baselines (seconds / epoch), where the components of PGDR show negligible effects on its efficiency. As shown in the Table 16, PGDR performs comparably to the baseline methods in terms of speed.
Appendix D Mean and Standard Deviation
For each experiment, we conduct three independent runs, with the reported values in the main experimental table representing the mean results. The corresponding standard deviations are listed in Table 17.
Method $q=0.1$ $q=0.2$
10-blurry 30-blurry 10-blurry 30-blurry
PiCO 0.24 0.32 0.19 0.30
+iCaRL 0.31 0.28 0.35 0.29
+BiC 0.78 0.81 0.95 0.79
+WA 0.41 0.52 0.38 0.36
+ER-ACE 0.36 0.47 0.39 0.56
+ANCL 0.69 0.92 1.06 0.70
+PGDR 0.88 0.89 0.94 0.78
Table 17: Standard deviation of methods on CIFAR100.
Appendix E Pseudo-code of PGDR